OpenStudy (anonymous):

help! Simplify. write without negative exponent. 1) (3x)^2(x)^-5 2)(-5x)^0 3) (-7x)^-2 4)(a^6b^5/10a^3)^3 5)(-20d^4/5b^3)^3

OpenStudy (blurbendy):

Well for the first one 3x^2 remains on top x^-5 is just 1/x^5 so, you're left with (3x)^2 / x^5 = 9/x^3 Do you understand all those steps?

OpenStudy (anonymous):

Yeah, sort of @blurbendy

OpenStudy (blurbendy):

what dont you get

OpenStudy (anonymous):

how did you get the x^3? @blurbendy

OpenStudy (blurbendy):

if you have something that's like x^2 / x^4 it's just 1 / x^2 the x'2s on top canceled with two of them on the bottom

OpenStudy (anonymous):

Can you help me with the rest of them? @blurbendy

OpenStudy (blurbendy):

the second one is easy. anything to the ^0 power is always 1

OpenStudy (blurbendy):

try number 3, what do you think?

OpenStudy (anonymous):

I have no clue, when I use a calculator I get a decimal

OpenStudy (anonymous):

@blurbendy

OpenStudy (blurbendy):

you dont need a calculator. just look at the first example. what happens if something is raised to a negative power?

OpenStudy (anonymous):

It's basically the opposite of the negative, so it will be positive? @blurbendy

OpenStudy (blurbendy):

the exponent will be. you end up with 1 / (-7x)^2 = 1 / 49x^2

OpenStudy (anonymous):

How would you do 4 and 5 they are fractions and have a few more exponents than the other ones? @blurbendy

OpenStudy (blurbendy):

you're left with a^3 on top, because 3 of them cancel with the a^3 on the bottom. then just raise everything to the third power (a^27b^15/1000)

OpenStudy (anonymous):

where did you get the 27 from? @blurbendy

OpenStudy (blurbendy):

(a^3)^3 = a^27

OpenStudy (anonymous):

Oh, okay. Now what about 5? @blurbendy

OpenStudy (blurbendy):

first divide 20 by 5 = 4 then cube everything = -64b^9d^12